# Tailored Forecasting from Short Time Series via Meta-learning

### v1

@@ III Discussion
In conclusion, METAFORS represents a significant step forward in data-driven forecasting of dynamical systems. By leveraging meta-learning to tailor forecasting models from related systems to new test systems, METAFORS mitigates key limitations of traditional memory-based models, enabling accurate predictions from limited data. The framework’s flexibility, efficiency, and generalization capabilities provide a robust foundation for tackling pressing forecasting challenges, from modeling climate dynamics to advancing real-time predictive analytics in fields such as public health and neuroscience.
IV Methods
IV.1 Forecasting with reservoir computers
@@
This section provides background for the description of our reservoir computing implementation of METAFORS to be presented in Section IV.2. The specific reservoir computing architecture that we employ is constructed after the one proposed by Jaeger and Haas in 2004Jaeger and Haas (2004). They referred to their version of a reservoir computer (RC) as an ‘echo state network’, but the two terms have become largely synonymous in many contexts. In their seminal work, Jaeger and Haas demonstrated the use of RCs in predicting and processing time series.
@@
The central component of an RC is a recurrent neural network, the reservoir. Each of its nodes, indexed by $i$, has an associated continuous-valued, time-dependent activation level $r_{i}(t)$. For a reservoir of size $N_{r}$ nodes, we refer to the vector ${\bm{r}(t)=[r_{1}(t),...,r_{N_{r}}(t)]^{T}}$, containing the activations of all of its nodes, as the reservoir state at time $t$. It evolves in response to an input signal according to a dynamical equation with a fixed discrete time-step, $\Delta t$:
@@
$+\lambda\tanh(A\bm{r}(t)+B\bm{u}(t)+C),$
where the $\tanh()$ function is applied element-wise, and the directed and weighted adjacency matrix $A$ is an ${N_{r}\times N_{r}}$ matrix specifying the strength and sign of interactions between each pair of reservoir nodes. We say that the reservoir has memory if ${\bm{r}(t+\Delta t)}$ depends on the reservoir past history, ${\bm{u}(t-m\Delta t)}$ for ${m>0}$, and (because, by Eq. 7, ${\bm{r}(t)}$ depends on ${\bm{u}(t-\Delta t)}$, ${\bm{r}(t-\Delta t)}$ depends on ${\bm{u}(t-2\Delta t)}$, and so on) this will be the case if the right hand side of Eq. 7 explicitly depends on ${\bm{r}(t)}$ (i.e., if ${(1-\lambda)}$ and/or the matrix $A$ are nonzero). The $N_{in}$-dimensional input ${\bm{u}(t)}$ is coupled to the reservoir nodes at time $t$ via an ${N_{r}\times N_{in}}$ input weight matrix $B$. $C$ is an $N_{r}$-dimensional random vector of biases, which serves to break symmetries in the dynamics of the reservoir nodes. $\lambda$, the leakage rate, influences the time-scale on which the reservoir state evolves; when the leakage rate is zero, ${\bm{r}(t+\Delta t)=\bm{r}(t)\ \forall\ t}$ and the reservoir does not evolve; when the leakage rate is one, the first term in Eq. 7 is zero and the reservoir ‘forgets’ its previous states more rapidly.
We use the rescompy python packageCanaday et al. (2024) to construct our RCs. The adjacency matrix, $A$, of each reservoir is a (sparse) random directed network with connection probability ${\langle d\rangle/N_{r}}$ for each pair of nodes, where $\langle d\rangle$ is the mean in-degree of the network. We assign non-zero elements of $A$ random values from a uniform distribution ${U[-1,1]}$. Then we normalize this randomly generated matrix such that its spectral radius (eigenvalue of largest absolute value) has some desired value, $\rho$. We generate a dense input matrix, $B$, and a bias vector, $C$, by choosing each entry from the uniform distributions ${U[-\sigma,\sigma]}$ and ${U[-\psi,\psi]}$, respectively. We refer to $\sigma$ as the input strength and to $\psi$ as the bias strength.
@@
$W_{out}$ is the ${N_{out}\times N_{r}}$ matrix that determines the linear combination. We call $W_{out}$ the reservoir’s output layer and refer to the combination of a reservoir (defined by its internal parameters: the adjacency matrix, $A$, input matrix, $B$, bias vector, $C$, and leakage rate, $\lambda$) and an output layer, $W_{out}$, as a reservoir computer (RC). When training an RC for forecasting tasks, we choose its output layer so that at every time-step over some training signal of $N_{train}$ evenly-spaced data points, i.e. with duration ${(N_{train}-1)\Delta t}$, the RC’s output closely matches its input at the next time-step:
@@
In this case, $N_{out}=N_{in}=N_{sys}$, where $N_{sys}$ is the number of observables we wish to predict. The internal parameters of the reservoir are unaltered during training. To calculate the output layer, we drive the reservoir with the training signal in the open-loop mode (Eqs. 7 and 2a) and minimize the ridge-regression cost function:
$\sum_{n=N_{trans}+1}^{N_{train}}\frac{\|W_{out}\bm{r}(n\Delta t)-\bm{u}(n\Delta t)\|^{2}}{N_{fit}}+\alpha_{F}\|W_{out}\|^{2},$ (10)
@@
Once we have trained a reservoir computer (RC) as above, we can use it to obtain predictions, $\hat{\bm{u}}(t)$, of $\bm{u}(t)$, for values of $t>t_{0}\geq N_{train}\Delta t$, where $t_{0}$ is the forecast start time. In the prediction phase, the RC evolves autonomously in closed-loop mode (Fig. 2B) by setting its input at each time-step to be its output from the previous time-step in a cyclic manner. The RC itself thus forms an autonomous dynamical system which evolves according to the coupled equations
@@
In general, the reservoir component of an RC that has been successfully trained to accomplish its time series prediction task satisfies the echo state propertyJaeger (2001); Lukoševičius (2012); Cucchi et al. (2022); Platt et al. (2021, 2022): if the reservoir is driven in the open-loop mode (Eqs. 7 and 2a) twice with the same input signal but different initial reservoir states, it will converge in both cases to the same trajectory after sufficient time. In other words, after some transient response of the reservoir has passed, its state becomes independent of its initial condition and depends only on the history of the input. One consequence of the echo state property is that any input trajectory ${\bm{u}}$ that evolves on some manifold, $M_{sys}$ will drive the reservoir to evolve along a trajectory, ${\bm{r}}$, such that once the transient period has passed, $\bm{r}$ and $\bm{u}$ are synchronized and $\bm{r}$ lies on some corresponding manifold, $M_{res}$, in the reservoir state-spaceLu, Hunt, and Ott (2018); Platt et al. (2021, 2022). The output layer of an RC trained on $\bm{u}$ maps points on or near $M_{res}$ to points on or near $M_{sys}$. For a fixed reservoir (defined by its internal parameters $A$, $B$, $C$, and $\lambda$), the structures of these manifolds and, consequently, of the RC’s output layer are determined solely by the dynamics of the system from which $\bm{u}$ is sampled – the initial conditions of the reservoir and of the time series $\bm{u}$ have negligible impact if $N_{trans}$ and $N_{train}$ are sufficiently large. We may thus build the following intuition, which is central to our RC implementation of METAFORS: an output layer, $W_{out}$, of an RC encodes the dynamics of the system on the manifold $M_{sys}$; the reservoir state, ${\bm{r}(t)}$, determines its phase.
@@
Typically, to initialize a forecast, we drive the reservoir in the open-loop mode (Eqs. 7 and 2a) with a sync signal, $\bm{u}_{sync}$, from the true system and then switch to the closed-loop mode (Eqs. 12b and 2b) to forecast from the end of $\bm{u}_{sync}$. The sync signal can be significantly shorter than would be sufficient to train an RC accurately. Hence, an RC that has been trained on a long time series to accurately capture the dynamics of $\bm{u}(t)$, can be used to predict from a different initial condition by starting from a comparatively short sync signal. However, if the sync signal is too short, the initial state error (1) will lead to an inaccurate forecast even if the output layer can accurately capture the dynamics of the true system.
IV.2 Our reservoir-computing implementation of METAFORS
Meta-learning for Tailored Forecasting from Related Time Series (METAFORS) leverages a library of long time series from related dynamical systems to build suitable forecasting models for similarly related test systems with limited available data. METAFORS uses two levels of learning to perform this task. Here, we use reservoir computing for both levels. In the first level, we train output layers for the ’forecaster reservoir’ separately on each available signal to construct a library of corresponding forecaster RCs. In the second level, the ’signal mapper’ RC, draws on all the dynamics represented in the library of forecaster RCs, and a short test signal from a system of interest, to both construct and initialize a suitable forecaster RC for that system. Our training scheme for the signal mapper RC has elements in common with schemes used in RC-based similarity learning that have been applied to image recognition and classificationKrishnagopal, Aloimonos, and Girvan (2018); Schaetti, Salomon, and Couturier (2016). In our case, the signal mapper infers similarity between short observational time series and learns a mapping from these short series to corresponding output layers (trainable parameters) and initial reservoir states (cold-start vectors) for a forecaster RC.
Because of their simplicity, accuracy, and efficiency in both short-term forecasting and long-term climate replication, we believe RCs are a particularly strong choice for building the forecaster used in the first level of learning. For the signal mapper in the second level of learning, we chose RCs primarily for their convenience, but they remain a robust and effective option. Nonetheless, we anticipate that other types of ML approaches could also perform well for these functions and might have advantages in certain situations. Alternative methods might offer comparable or improved performance for the forecaster, additional flexibility or interpretability for the signal mapper, or even complementary benefits that enhance the overall framework’s adaptability to different scenarios.
IV.2.1 Requirements and scope
METAFORS requires that we have available a short observed test signal or cue, $\bm{s}_{test}$, from the dynamical system we wish to predict as well as a collection, or library, $\{\bm{L}_{i}\},\>{i=1,...,N_{L}}$, of $N_{L}$ long signals from systems that exhibit similar dynamics to the test system. In our RC implementation of METAFORS, the data from each time series must be sequential and sampled at even intervals, $\Delta t$, and must also be of the same dimension, which we denote $N_{sys}$. We emphasize that the library and test signals need not contain full information of the system state at each time step. $N_{sys}$ is merely the number of observables we have data for and wish to predict. We do not require that all long signals have equal duration, but we assume that each is sufficiently long to train an RC well. METAFORS is most useful when the duration of the test signal, ${t_{test}=(N_{test}-1)\Delta t}$, where $N_{test}$ is the number of data points in the test signal, is insufficient to train an RC directly. METAFORS’ ability to cold-start the forecaster RC is most useful when $t_{test}$ is insufficient to synchronize its reservoir state to the test signal.
IV.2.2 Constructing the library
  1. (2a)
  2. (2b)
    $\bm{s}_{ij}(k\Delta t)=\bm{L}_{i}(j\Delta t+k\Delta t)\\ \ \forall\ \begin{cases}1\leq i\leq N_{L}\\ j>=N_{trans},\\ 0\leq k\leq N_{test}-1\end{cases}$ (13)
    where $\bm{s}_{ij}$ denotes the $j^{th}$ sub-signal extracted from long signal $\bm{L}_{i}$ and $N_{trans}$ is the transient time, as in Eq. 10. Here $j$ also indexes the time step at the start of the short signal, since we utilize the full set of possible short signals after the transient. Although we note that METAFORS can still perform strongly when short signals are subsampled from the long library signals.
  3. (2c)
    $\bm{r}_{ij}(0)=\bm{r}_{i}(j\Delta t)$ (14)
    from the stored reservoir trajectories, $\{\bm{r}_{i}\}$. ${\bm{r}_{ij}(0)}$ is a constructed cold-start vector, $\bm{m}_{ij}$, for the test signal $\bm{s}_{ij}$ (Fig. 1B).
Each library member in METAFORS comprises a long signal and its corresponding trained forecaster model. The set of short signals, associated initial reservoir states, and trained output layers ${\{(\bm{s}_{ij},\bm{r}_{ij}(0),W_{out}^{i})\}}$ forms the training data for the signal mapper.
IV.2.3 Training the signal mapper RC
  1. (3a)
  2. (3b)
    $W_{SM}=PR^{T}\left(RR^{T}+\alpha_{SM}N_{short}I\right)^{-1},$ (15)
    where $N_{short}$ is the number of short signals in the library (the number of training pairs), and $R$ (${N_{SM}\times N_{short}}$) and $P$ (${(N_{sys}+1)N_{F}\times N_{short}}$) are the horizontal concatenations of all final signal mapper states and all target pairs, ${\bm{p}_{ij}=[\bm{r}_{ij}(0),W_{flat}^{i}]^{T}}$, respectively. $W_{flat}^{i}$ is a vector representation of the output-layer $W_{out}^{i}$ with dimension $N_{F}N_{sys}$.
IV.2.4 Making predictions
  1. (4a)
  2. (4b)
    We combine the inferred output layer, $W_{out}^{test}$, with the forecaster reservoir to construct a tailored forecaster RC for the test system. We then synchronize this RC to the test signal in open-loop mode (Fig. 2A) starting from the initial state ${\bm{r}(0)=\bm{r}_{test}(0)}$, and close the loop (Fig. 2B) after time ${t_{test}=(N_{test}-1)\Delta t}$ to forecast from the end of the test signal. The reservoir inputs are then:
    Thus, ${\hat{\bm{u}}(N_{test}\Delta t)}$ is the first output of the forecaster RC to be fed back as its input. In other words, $t_{test}$ corresponds to the forecast start time $t_{0}$ in Section IV.1.
IV.3 Testbeds for probing the utility of METAFORS
IV.3.1 The logistic and Gauss iterated maps
@@
For each long signal in the library, we simulate the logistic/Gauss iterated map for $2000$ total iterations. We discard the first $1000$ iterations of each to ensure that the time series used for training are well-converged to the system’s attractor. Each signal’s remaining ${N_{train}=1000}$ points constitute a long library signal. We use the next ${N_{trans}=50}$ iterations of each signal as a transient period to synchronize the internal state of the forecaster RC to the input series, and then fit the trainable parameters of the forecaster, contained in its output layer, to the remaining $N_{fit}=950$ iterations. We forecast $N_{for}=1000$ iterations beyond the end of each test signal. Since, in these experiments, we are interested only in the long-term statistics, or climate, of a forecast, we discard the first $500$ of these predicted points before calculating cumulative probability distributions or plotting bifurcation diagrams.
As shown in Table 1, the reservoir hyperparameters that we use for our experiments with the logistic map only (Section II.3) and for our experiments with both the logistic and Gauss iterated maps (Section II.4) are identical except for the input strengths, $\sigma_{F}$ and $\sigma_{SM}$. We chose the size, mean in-degree, and bias of both networks to have values that typically allow for reasonably accurate forecasting with reservoir computers, and performed no experiment-specific tuning of these values. We chose the leakage rate and spectral radius of the signal mapper such that its memory is long. This long memory ensures that the final reservoir state of the signal mapper, after it has received a test signal as input, is influenced by many, or all, of the data points in the test signal. We chose the input strength of both the forecaster and signal mapper networks such that the average of the product of the input weight matrix and the input data from the library is approximately one, ${\langle B\bm{u}_{i}(t)\rangle_{i,t}=\langle B\bm{L}_{i}(t)\rangle_{i,t}\approx 1}$. We chose the remaining hyperparameters – the regularization strengths, forecaster leakage rate, and forecaster spectral radius – by coarse hand-tuning to allow for good, but not necessarily optimal performance. While more robust hyperparameter tuning may improve performance overall, our priority is to compare the relative performances of simple baseline methods on familiar systems, rather than to obtain highly optimized forecasts.

### v2

@@ Abstract
In conclusion, METAFORS represents a significant advance in data-driven forecasting of dynamical systems. By leveraging meta-learning to construct tailored forecasting models, it mitigates key limitations of traditional memory-based approaches, enabling accurate predictions from limited data. Its flexibility, efficiency, and generalization capabilities make it a powerful tool for tackling pressing forecasting challenges across fields such as climate science, neuroscienceDe Matola and Miniussi (2025), and public health.
BrickRedIVMethods
In this section, we build on the earlier implementation overview to more thoroughly detail our reservoir computing implementation of METAFORS. The supplementary material contains additional background information on forecasting time series with reservoir computers (Section˜.1), and details of our experiment setups (Section˜.2) and baseline comparison methods (Section˜.3).
IV.1Our reservoir-computing implementation of METAFORS
METAFORS uses two levels of learning to build and cold-start tailored forecasting models for data-limited dynamical systems by leveraging a library of models trained on potentially related long time series. Here, we use reservoir computers (RCs) for both learning levels. In the first level, we construct a library of forecaster RCs by training a different output layer for the ‘forecaster reservoir’ on each available long signal. In the second level, a ‘signal mapper’ RC, draws on all the dynamics represented in the library of forecaster RCs, and a short cue signal to both construct and initialize a suitable forecaster RC for that cue. Our training scheme for the signal mapper RC has elements in common with schemes used in RC-based similarity learning that have been applied to image recognition and classificationKrishnagopal, Aloimonos, and Girvan (2018); Schaetti, Salomon, and Couturier (2016). In our case, the signal mapper infers similarity between short observed time series and learns a mapping from these short series to corresponding output layers (trainable parameters) and initial reservoir states (cold-start vectors).
We believe that RCs are a particularly strong choice for building the forecaster used in the first level of learning because of their simplicity, accuracy, and efficiency in both short-term forecasting and long-term climate replication. For the signal mapper in the second level of learning, however, we chose RCs primarily for their convenience. They remain a robust and effective option, but we nonetheless anticipate that other types of ML approaches could also perform effectively and might be advantageous in certain situations. Alternative methods might offer comparable or improved performance for the forecaster, additional flexibility or interpretability for the signal mapper, or even complementary benefits that enhance the overall framework’s adaptability to different scenarios.
IV.1.1Requirements and scope
METAFORS requires that we have available a short observed signal or cue, $\bm{s}_{new}$ (denoted $\bm{s}_{test}$ in the case of our experiments with simple test systems), from the dynamical system we wish to predict as well as a collection, or library, $\{\bm{L}_{i}\},\>{i=1,...,N_{L}}$, of $N_{L}$ long signals from systems that exhibit similar dynamics to the test system. In our RC implementation of METAFORS, the data from each time series must be sequential and sampled at even intervals, $\Delta t$, and must also be of the same dimension, which we denote $N_{sys}$. We emphasize that the library and test signals need not contain full information of the system state at each time step. $N_{sys}$ is merely the number of observables we wish to predict. We do not require that all long signals have equal duration, but we assume that each is sufficiently long to train an RC well. METAFORS is most useful when the duration of the test signal, ${t_{test}=(N_{test}-1)\Delta t}$, where $N_{test}$ is the number of data points in the signal, is insufficient to train an RC directly. METAFORS’ ability to cold-start the forecaster RC is most useful when $t_{test}$ is insufficient to synchronize the forecaster reservoir state to the test signal.
IV.1.2Constructing the library
  1. (2a)
  2. (2b)
    $\bm{s}_{ij}(k\Delta t)=\bm{L}_{i}(j\Delta t+k\Delta t)\\ \ \forall\ \begin{cases}1\leq i\leq N_{L}\\ j\geq N_{trans},\\ 0\leq k\leq N_{test}-1\end{cases}$ (6)
    where $\bm{s}_{ij}$ denotes the $j^{th}$ sub-signal extracted from long signal $\bm{L}_{i}$ and $N_{trans}$ is the transient time, as in Eq. 12. Here $j$ also indexes the time step at the start of the short signal since we utilize all possible short signals after the transient, although we note that METAFORS can still perform strongly when short signals are subsampled from the long library signals.
  3. (2c)
    $\bm{r}_{ij}(0)=\bm{r}_{i}(j\Delta t)$ (7)
    from the stored reservoir trajectories, $\{\bm{r}_{i}\}$. ${\bm{r}_{ij}(0)}$ is a constructed cold-start vector, $\bm{m}_{ij}$, for the signal $\bm{s}_{ij}$ (Fig. 1B).
Each library member in METAFORS comprises a long signal and its corresponding trained forecaster model. The set of triplets of short signals, associated initial reservoir states, and trained output layers ${\{(\bm{s}_{ij},\bm{r}_{ij}(0),W_{out}^{i})\}}$ forms the training data for the signal mapper.
IV.1.3Training the signal mapper RC
  1. (3a)
  2. (3b)
    $W_{SM}=PR^{T}\left(RR^{T}+\alpha_{SM}N_{short}I\right)^{-1},$ (8)
    where $N_{short}$ is the number of short signals in the library (the number of training pairs), and $R$ (${N_{SM}\times N_{short}}$) and $P$ ${\left((N_{sys}+1)N_{F}\times N_{short}\right)}$ are the horizontal concatenations of all final signal mapper states and all target pairs, ${\bm{p}_{ij}=[\bm{r}_{ij}(0),W_{flat}^{i}]^{T}}$, respectively. $W_{flat}^{i}$ is an $N_{F}N_{sys}$-dimensional vector representation of output layer $W_{out}^{i}$.
IV.1.4Making predictions
  1. (4a)
  2. (4b)
    We construct a tailored forecaster RC for the test system by combining the inferred output layer, $W_{out}^{test}$, with the forecaster reservoir. We then synchronize this RC to the test signal in open-loop mode (Fig. 2A) starting from the initial state ${\bm{r}(0)=\bm{r}_{test}(0)}$, and close the loop (Fig. 2B) after time ${t_{test}=(N_{test}-1)\Delta t}$ to forecast from the end of the test signal. The reservoir inputs are then:
    Thus, ${\hat{\bm{u}}(N_{test}\Delta t)}$ is the first output of the forecaster RC to be fed back as its input. In other words, $t_{test}$ corresponds to the forecast start time $t_{0}$ in Section˜.1.
BrickRedVReferences
Acknowledgements.
We thank Daniel Canaday for helpful conversations, insights, and suggestions. We also acknowledge the University of Maryland supercomputing resources (http://hpcc.umd.edu) made available for conducting the research reported in this paper. This work was supported in part by DARPA under contract W31P4Q-20-C-0077. The contributions of D.N., of M.G., and of E.O. and B.H. were also supported, respectively, by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 1840340, by ONR Grant No. N000142212656, and by ONR Grant No.N00014-22-1-2319. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation, the Department of Defense, or the U.S. Government.
BrickRedVICode Availability:
Codes used to run the numerical experiments performed for this paper are available at the following public repository: https://github.com/nortondeclan/METAFORS.
BrickRedVIISupplementary Material
.1 Forecasting with reservoir computers
@@ Acknowledgements.
The specific reservoir computing architecture that we employ is constructed after the one proposed by Jaeger and Haas in 2004Jaeger and Haas (2004). They referred to their version of a reservoir computer (RC) as an ‘echo state network,’ but the two terms have been used interchangeably in many contexts. In their seminal work, Jaeger and Haas demonstrated the use of RCs in predicting and processing time series.
@@
The central component of an RC is a recurrent neural network, the reservoir. Each of its nodes, indexed by $i$, has an associated continuous-valued, time-dependent activation level $r_{i}(t)$. For a reservoir of size $N_{r}$ nodes, we refer to the vector ${\bm{r}(t)=[r_{1}(t),...,r_{N_{r}}(t)]^{T}}$, containing the activations of all of its nodes, as the reservoir state at time $t$. It evolves in response to an input signal according to a dynamical equation with a fixed discrete time step, $\Delta t$:
@@
$+\lambda\tanh(A\bm{r}(t)+B\bm{u}(t)+\bm{c}),$
where the $\tanh()$ function is applied element-wise, and the directed and weighted ${N_{r}\times N_{r}}$ adjacency matrix $A$ specifies the strength and sign of interactions between each pair of reservoir nodes. The $N_{in}$-dimensional input ${\bm{u}(t)}$ is coupled to the reservoir nodes at time $t$ via an ${N_{r}\times N_{in}}$ input weight matrix $B$. An $N_{r}$-dimensional random vector of biases, $\bm{c}$, serves to break symmetries in the dynamics of the reservoir nodes. We say that the reservoir has memory if ${\bm{r}(t+\Delta t)}$ depends on the reservoir past history of inputs, ${\bm{u}(t-m\Delta t)}$ for ${m>0}$, and (because, by Eq. 9, ${\bm{r}(t)}$ depends on ${\bm{u}(t-\Delta t)}$, ${\bm{r}(t-\Delta t)}$ depends on ${\bm{u}(t-2\Delta t)}$, and so on) this will be the case if the right hand side of Eq. 9 explicitly depends on ${\bm{r}(t)}$ (i.e., if ${(1-\lambda)}$ and/or the matrix $A$ are nonzero). The leakage rate, $\lambda$, thus influences the time-scale on which the reservoir state evolves; when the leakage rate is zero, ${\bm{r}(t+\Delta t)=\bm{r}(t)\ \forall\ t}$ and the reservoir does not evolve; when the leakage rate is one, the first term in Eq. 9 is zero and the reservoir ‘forgets’ its previous states more rapidly.
We use the rescompy python packageCanaday et al. (2024) to construct our RCs. The adjacency matrix, $A$, of each reservoir is a (sparse) random directed network with connection probability ${\langle d\rangle/N_{r}}$ for each pair of nodes, where $\langle d\rangle$ is the mean in-degree of the network. We assign non-zero elements of $A$ random values from a uniform distribution ${U[-1,1]}$. Then we normalize this randomly generated matrix such that its spectral radius (eigenvalue of largest absolute value) has some desired value, $\rho$. We generate a dense input matrix, $B$, and a bias vector, $\bm{c}$, by choosing each entry from the uniform distributions ${U[-\sigma,\sigma]}$ and ${U[-\psi,\psi]}$, respectively. We refer to $\sigma$ as the input strength and to $\psi$ as the bias strength.
@@
$W_{out}$ is the ${N_{out}\times N_{r}}$ matrix that determines the linear combination. We call $W_{out}$ the reservoir’s output layer and refer to the combination of a reservoir (defined by its internal parameters: the adjacency matrix, $A$, input matrix, $B$, bias vector, $\bm{c}$, and leakage rate, $\lambda$) and an output layer, $W_{out}$, as a reservoir computer (RC). When training an RC for forecasting tasks, we choose its output layer so that at every time step over some training signal of $N_{train}$ evenly-spaced data points, i.e. with duration ${(N_{train}-1)\Delta t}$, the RC’s output closely matches its input at the next time step:
@@
In this case, $N_{out}=N_{in}=N_{sys}$, where $N_{sys}$ is the number of observables we wish to predict. The internal parameters of the reservoir ($A$, $B$, $\bm{c}$, and $\lambda$) are not altered during training, or afterwards. To calculate the output layer, we drive the reservoir with the training signal in the open-loop mode (Eq. 9 and Fig. 2A) and minimize the ridge-regression cost function:
$\sum_{n=N_{trans}}^{N_{train}-1}\frac{\|W_{out}\bm{r}(n\Delta t)-\bm{u}(n\Delta t)\|^{2}}{N_{fit}}+\alpha_{F}\|W_{out}\|^{2},$ (12)
@@
Once we have trained a reservoir computer (RC) as above, we can use it to obtain predictions, $\hat{\bm{u}}(t)$, of $\bm{u}(t)$, for values of $t>t_{0}\geq(N_{train}-1)\Delta t$, where $t_{0}$ is the forecast start time. In the prediction phase, the RC evolves autonomously in closed-loop mode (Fig. 2B) by setting its input at each time step to be its output from the previous time step in a cyclic manner. The RC itself thus forms an autonomous dynamical system which evolves according to the coupled equations
@@
In general, the reservoir component of an RC that has been successfully trained to accomplish its time series prediction task satisfies the echo state propertyJaeger (2001); Lukoševičius (2012); Cucchi et al. (2022); Platt et al. (2021, 2022): if the reservoir is driven in the open-loop mode (Eq. 9 and Fig. 2A) twice with the same input signal but different initial reservoir states, it will converge in both cases to the same trajectory after sufficient time. In other words, after some transient response of the reservoir has passed, its state becomes independent of its initial condition and depends only on the history of the input. One consequence of the echo state property is that any input trajectory ${\bm{u}}$ that evolves on some manifold, $M_{sys}$ will drive the reservoir to evolve along a trajectory, ${\bm{r}}$, such that once the transient period has passed, $\bm{r}$ and $\bm{u}$ are synchronized and $\bm{r}$ lies on some corresponding manifold, $M_{res}$, in the reservoir state-spaceLu, Hunt, and Ott (2018); Platt et al. (2021, 2022). The output layer of an RC trained on $\bm{u}$ maps points on or near $M_{res}$ to points on or near $M_{sys}$. For a fixed reservoir (defined by its internal parameters $A$, $B$, $\bm{c}$, and $\lambda$), the structures of these manifolds and, consequently, of the RC’s output layer are determined solely by the dynamics of the system from which $\bm{u}$ is sampled – the initial conditions of the reservoir and of the time series $\bm{u}$ have negligible impact if $N_{trans}$ and $N_{train}$ are sufficiently large. We may thus build the following intuition, which is central to our RC implementation of METAFORS: an output layer, $W_{out}$, of an RC encodes the dynamics of the system on the manifold $M_{sys}$; the reservoir state, ${\bm{r}(t)}$, determines its phase.
@@
Typically, to initialize a forecast, we drive the reservoir in the open-loop mode (Eq. 9 and Fig. 2A) with a sync signal, $\bm{u}_{sync}$, from the true system and then switch to the closed-loop mode (Eq. 14b and Fig. 2B) to forecast from the end of $\bm{u}_{sync}$. The sync signal can be significantly shorter than would be sufficient to train an RC accurately. Hence, an RC that has been trained on a long time series to accurately capture the dynamics of $\bm{u}(t)$, can be used to predict from a different initial condition by starting from a comparatively short sync signal. However, if the sync signal is too short, the initial state error (1) will lead to an inaccurate forecast even if the output layer can accurately capture the dynamics of the true system.
.2 Testbeds for probing the utility of METAFORS
.2.1 The logistic and Gauss iterated maps
We exclude periodic trajectories from the training library in our experiments with the logistic and Gauss iterated maps because these trajectories poorly sample the state space of the maps’ dynamics. A non-parametric forecasting model, such as an RC, trained on an orbit with a short periodicity length may not learn a sufficiently complex representation of the dynamics. As a result, its inclusion in the training library can hamper generalization.
@@
For each long signal in the library, we simulate the logistic/Gauss iterated map for 200020002000 total iterations. We discard the first 100010001000 iterations of each to ensure that the time series used for training are well-converged to the system’s attractor. Each signal’s remaining ${N_{train}=1000}$ points constitute a long library signal. We use the next ${N_{trans}=50}$ iterations of each signal as a transient period to synchronize the internal state of the forecaster RC to the input series, and then fit the trainable parameters of the forecaster, contained in its output layer, to the remaining $N_{fit}=950$ iterations. We forecast $N_{for}=1000$ iterations beyond the end of each test signal. Since, in these experiments, we are interested only in the long-term statistics, or climate, of a forecast, we discard the first 500500500 of these predicted points before calculating cumulative probability distributions or plotting bifurcation diagrams.